# Scientific judgment drifts over time in AI ideation

### v1

@@ Discussion
Evaluating scientific ideas is uniquely challenging. It requires balancing originality, feasibility, potential impact, and many other value dimensions that are inherently uncertain and context-dependent. Our longitudinal study shows that these human evaluative criteria are not fixed in time. Ratings for an unchanged control idea drifted upward between waves, and test–retest reliability across dimensions was only moderate and heterogeneous. When we tuned an AI evaluator to match a one-time snapshot of human criteria (Wave 1) and used it to select ideas for Wave 2, the apparent gains vanished once we corrected for concurrent drift using the repeated control. Optimizing to a fixed snapshot improved a transient metric but did not yield genuine and persistent improvements in human ratings.
@@
At the same time, we found that the “structure” of human judgment remained largely stable. The relative importance scientists assigned to different dimensions, such as originality, clarify, or implementability, did not change significantly across waves. This suggests that while overall leniency and absolute ratings drifted, the hierarchy of values that underlies scientific idea evaluation persisted. Human evaluation, therefore, is not unreliable noise, but a dynamical system with stable internal structure and shifting baseline calibration. This distinction matters for how we design and interpret human-in-the-loop AI ideation systems. Models that assume stability risk overfitting to transient conditions, but models that explicitly track and adapt to evolving human criteria may bring genuine and persistent advancements over time.
This idea aligns with lessons from adjacent fields. In recommender systems, temporal drift in user preferences is expected and routinely modeled (?, ?). In behavioral and cognitive research, intra-rater variability is treated as an intrinsic property of human judgment rather than an error to be eliminated (?, ?). Scientific ideation systems can adopt similar strategies by incorporating temporal anchors, repeated controls, and adaptive evaluators.
@@
Our work represents an initial step. There are several future opportunities to improve the study. Our two-wave design demonstrates the existence of preference drift but does not map its longer-term dynamics. Future work spanning more waves could reveal whether evaluation criteria eventually stabilize, oscillate, or continue to drift over a longer period of time. Our human evaluative model was a simple linear combination of dimensions; more expressive models such as dynamic latent-variable approaches may capture additional nuance of the human criteria such as nonlinear or interaction effects.
We measured perceived scientific quality, but not yet downstream scientific impact. Understanding whether early-stage idea ratings predict discoveries, publications, or funded projects will require long-term field experiments and future advancements in automated scientific experiments. Though our study covers multiple departments and different level of researchers, our sample comprised scientists from a single institution. Cross-community replication could reveal whether drift patterns generalize across scientific cultures. Moreover, our results are based on a representative yet specific LLM-based ideation system and alignment method; future studies could explore how these findings generalize across different models and system designs. With more accumulated data from the community, more advanced alignment methods such as model fine-tuning could be implemented. Additionally, due to the consideration of obtaining higher response rates, we only included one repeated control idea per participant; more controls could further reduce variance. On the other hand, while repeated items is a standard way of measuring test-retest reliability, we cannot exclude familiarity effects as confounding variables that might have influenced ratings change. Future work could explore more sophisticated designs to address these risks.
@@ Conclusion
Human evaluation of scientific ideas is not fixed. In a two-wave study, we observed measurable temporal drift in overall quality ratings of the same ideas but largely stable internal structure in how scientists weigh different value dimensions. Aligning an AI ideation system to a one-time snapshot of human judgment produced transient gains that vanished once drift was considered. These results call for longitudinal and drift-aware evaluation protocols that incorporate repeated controls, report stability metrics, and explicitly model the dynamics of human judgment. Recognizing that scientific value judgments are dynamic rather than static may be key to building AI systems that not only generate ideas, but continue to evolve with the scientists who assess, improve, and perhaps take inspirations from them.
@@ Acknowledgments
[…]
@@ 1 Participants
Participants were recruited from the following six departments at Duke University: Biomedical Engineering, Computer Science, Electrical and Computer Engineering, Mechanical Engineering and Material Science, Psychology and Neuroscience, and Sociology. We filtered for active researchers to invite to our study. The criteria was having published at least one first- or last-author paper since January 2020 that has been cited. In the first wave of our study, we received 41 responses from faculty members and 54 responses from non-faculty researchers. Wave two surveys were sent to these 95 respondents. Researchers who responded to both waves were included in our final study. The Faculty group contained 21 participants; the postdoctoral fellow, graduate students, and research staff group contained 39 participants. All non-faculty participants’ advisors are included in the faculty group. Since non-faculty participants often don’t have papers that meet our selection criteria. However, we assume they are familiar with their advisors’ research, hence we assign the same survey content of the faculty member to their group members.
@@ 2 Idea evaluation survey
For each wave, each participant completed a customized Google Forms survey lasting approximately 20 to 30 minutes. In each survey, participants evaluated seven ideas: six AI-generated ideas and one repeated human-written control idea (presented in random order). For each idea, participants rated eight scientific value dimensions: Originality, Paradigm relatedness, Acceptability, Implementability, Applicability, Effectiveness, Implicational explicitness, and Clarity, along with an Overall Quality score (0-10 scale). Detailed instructions and prompts are provided in SM 13.
@@ 3 Structured idea representations
When extracting ideas from existing papers, the pdf file was first converted to text using pymupdf4llm (?). Then the idea was extracted using an LLM model (gpt-40-mini) with the prompt provided in SM 11. The temperature was set to 0.2 and top_p to 0.1. For both extracting and generating ideas, we used OpenAI’s structured output API to ensure the desired format is enforced.
@@ 4.1.2 Domain-compatibility pre-screening
Because not all cross-department pairings were equally promising, we performed a preliminary compatibility screen. For each department pair (e.g. [Biomedical Engineer, Sociology]), we sampled one conventional paper and ten candidate novel papers. For each of the ten pairs, we prompted the generator five times, yielding 50 ideas. We then scored these idea with our heuristic LLM evaluator (SM 4.2). The final compatibility of a pair of departments was calculated by the average LLM evaluator rating. Compatibility matrix are shown in Fig. S1
@@ 4.1.3 Model and prompting
We used OpenAI’s gpt-4o-2024-08-06 as the generator. For idea generation we kept temperature and top_p at their defaults (1.0) to encourage diversity. We used OpenAi’s structured output API to ensure that ideas were returned in a standardized structure SM 3. To avoid trivial re-statements of known work, the prompt explicitly instructed the model to be transformative, concrete, and not similar to existing works. The full generation prompt can be found in SM 12.
@@ 4.2 Evaluator design
Each generated idea was evaluated by an automated LLM-based evaluator that rated the same eight dimensions as human participants. To ground evaluations in existing literature, we implemented a retrieval-augmented pipeline. A gpt-4o-mini model generated five keyword summaries per idea, which were submitted to Semantic Scholars with five such queries. The retrieved papers’ titles, tldrs, and abstracts were saved into a pool of potentially relevant papers. We then filtered this pool of papers by their embedding similarity to the idea to be evaluated and kept those with a similarity higher than 0.88. We used OpenAI’s text-embedding-large model for this step. The threshold was decided after manual inspection for relevancy on a small set of examples.
The evaluation prompt (SM 13) used the same instructions and dimension anchors as those shown to human raters. For more consistent evaluation, we follow best practices in using LLMs as evaluators and compute a probability-weighted summation of the output token scores as the final score (?, ?, ?).
@@ 6 Predictors of rating change
Time gap. Since the time gap distribution was skewed (see SM 9), we took the log value of the days between responses.
@@ 13 Evaluation Instructions/Prompt
Originality 0.537 0.322 0.693
Paradigm relatedness 0.508 0.313 0.653
Acceptability 0.636 0.407 0.782
Implementability 0.570 0.258 0.791
Applicability 0.486 0.189 0.711
Effectiveness 0.532 0.282 0.705
Implicational explicitness 0.372 0.098 0.572
Clarity 0.396 0.115 0.618
@@
Originality 0.0 0.0 0.0 0.382 0.611 False
Paradigm relatedness 0.0 0.0 0.0 0.296 0.611 False
Acceptability 0.0 0.0 0.0 0.827 0.845 False
Implementability 0.0 0.0 0.0 0.326 0.611 False
Applicability 0.0 0.0 0.0 0.845 0.845 False
Effectiveness 0.0 0.0 0.0 0.226 0.611 False
Implicational explicitness 0.0 0.0 0.0 0.257 0.611 False
Clarity 0.0 0.0 0.0 0.695 0.845 False
Table S3: Mixed model on control Overall (drivers of change; z-scored covariates)
@@
Intercept 5.984 0.362 16.546 0.000 5.275 6.693
Expertise 0.149 0.458 0.327 0.744 0.747-0.747 1.046
Wave 0.449 0.163 2.752 0.006 0.129 0.769
Position 0.103 0.101 1.020 0.308 0.095-0.095 0.301
Context 0.143 0.202 0.710 0.478 0.253-0.253 0.540
Time Gap 0.199-0.199 0.229 0.868-0.868 0.385 0.649-0.649 0.250
Time of Day 0.052-0.052 0.115 0.453-0.453 0.651 0.278-0.278 0.174
S-wave 1.727 0.165 10.488 0.000 1.404 2.050

### v2

@@ Discussion
Evaluating scientific ideas is uniquely challenging. It requires balancing originality, feasibility, potential impact, and many other value dimensions that are inherently uncertain and context-dependent. Our longitudinal study shows that these human evaluative criteria are not fixed in time. Ratings for an unchanged control idea showed a small positive mean trend between waves, and test–retest reliability across dimensions was only moderate and heterogeneous indicating that repeated judgments of the same idea can vary substantially even within the same participant. When we tuned an AI evaluator to match a one-time snapshot of human criteria (Wave 1) and used it to select ideas for Wave 2, the apparent gains vanished once we corrected for concurrent drift using the repeated control. Optimizing to a fixed snapshot improved a transient metric but did not yield genuine and persistent improvements in human ratings.
@@
At the same time, we found that the “structure” of human judgment remained largely stable. The relative importance scientists assigned to different dimensions, such as originality, clarity, or implementability, did not change significantly across waves. This suggests that while overall leniency and absolute ratings drifted, the hierarchy of values that underlies scientific idea evaluation persisted. Human evaluation, therefore, is not unreliable noise, but a dynamical system with stable internal structure and shifting baseline calibration. This distinction matters for how we design and interpret human-in-the-loop AI ideation systems. Models that assume stability risk overfitting to transient conditions, but models that explicitly track and adapt to evolving human criteria may bring genuine and persistent advancements over time.
The mechanisms underlying these changes remain an open question. Our analysis finds no evidence that within-session order effects, local contextual comparisons, time-of-day, or time gap explain the observed variability. Several other explanations remain plausible. Procedural familiarity, cross-wave contrast/adaptation, and deeper shifts in scientific values driven by exposure to newer information or an evolving field could be contributors. Disentangling these mechanisms will require richer experimental designs. For example, a multi-arm Wave 2 design including aligned, unaligned, and deliberately “reverse-aligned” ideas could more directly test whether improvements persist beyond concurrent baseline changes. Such designs must balance causal resolution with survey burden and potential fatigue effects.
Our finding aligns with lessons from adjacent fields. In recommender systems, temporal drift in user preferences is expected and routinely modeled (?, ?). In behavioral and cognitive research, intra-rater variability is treated as an intrinsic property of human judgment rather than an error to be eliminated (?, ?). Scientific ideation systems can adopt similar strategies by incorporating temporal anchors, repeated controls, and adaptive evaluators.
@@
Our work represents an initial step. There are several future opportunities to improve the study. Our two-wave design demonstrates the existence of preference drift but does not map its longer-term dynamics. Future work spanning more waves could reveal whether evaluation criteria eventually stabilize, oscillate, or continue to drift over a longer period of time.
Our human evaluative model was a simple linear combination of dimensions; more expressive models such as dynamic latent-variable approaches may capture additional nuance of the human criteria such as nonlinear or interaction effects. We measured perceived scientific quality, but not yet downstream scientific impact. Understanding whether early-stage idea ratings predict discoveries, publications, or funded projects will require long-term field experiments and future advancements in automated scientific experiments.
Our study uses one repeated control idea per participant. While the 63 distinct control ideas provide robustness at the population level, multiple repeated controls per participant would improve estimation of within-person variability and help further disentangle procedural familiarity from substantive criteria change. We made a pragmatic tradeoff here: additional repeated items would lengthen an already 20-30 minute survey and could introduce fatigue effects. Future studies could consider shortening idea formats to keep survey length manageable.
Finally, though our study covers multiple departments and different level of researchers, our sample comprised scientists from a single institution. Cross-community replication could reveal whether drift patterns generalize across scientific cultures. Moreover, our results are based on a representative yet specific LLM-based ideation system and alignment method; future studies could explore how these findings generalize across different models and system designs. With more accumulated data from the community, more advanced alignment methods such as model fine-tuning could be implemented. Additionally, due to the consideration of obtaining higher response rates, we only included one repeated control idea per participant; more controls could further reduce variance. On the other hand, while repeated items is a standard way of measuring test-retest reliability, we cannot exclude familiarity effects as confounding variables that might have influenced ratings change. Future work could explore more sophisticated designs to address these risks.
@@ Conclusion
Human evaluation of scientific ideas is not fixed. In a two-wave study, we find that repeated ratings of the same ideas exhibit only moderate reliability, alongside small but variable changes in overall scores, while the relative importance assigned to different evaluation dimensions remains stable. Aligning an AI ideation system to a one-time snapshot of human judgment produced transient gains that vanished once drift was considered. These results call for longitudinal and drift-aware evaluation protocols that incorporate repeated controls, report stability metrics, and explicitly model the dynamics of human judgment. Recognizing that scientific value judgments are dynamic rather than static may be key to building AI systems that not only generate ideas, but continue to evolve with the scientists who assess, improve, and perhaps take inspiration from them.
@@ Acknowledgments
Lingyu Zhang1, Mitchell Wang2, Boyuan Chen1,2,3∗
@@ 1 Participants
Participants were recruited from the following six departments at Duke University: Biomedical Engineering, Computer Science, Electrical and Computer Engineering, Mechanical Engineering and Materials Science, Psychology and Neuroscience, and Sociology.
Faculty (seed) recruitment and filtering. To identify active faculty researchers and to obtain a publication set for customizing survey topics, we filtered faculty candidates by whether they had at least one first- or last-author paper published since January 2020 with external citations. This filter was used to target active labs and to support publication-based customization; it was not intended as a hard eligibility constraint for non-faculty participants.
Non-faculty recruitment and rationale. Non-faculty participants (postdoctoral fellows, graduate students, and research staff) were recruited from the corresponding faculty participants’ labs or field-adjacent labs as active researchers in the same topical area. Because non-faculty researchers may not yet have first/last-author publications meeting the faculty filter, we did not apply that criterion to non-faculty participation. To avoid introducing additional topic variability, non-faculty participants evaluated the same domain-customized survey content as their advisor or most matching faculty participant.
For the first wave of our study, we received 41 responses from faculty members. We used these response ratings to conduct the calibration step that yielded the Wave 2 candidate idea pools. Wave 2 surveys are only sent to researchers who responded to the Wave 1 survey. Researchers who responded to both waves were our final participants, comprising 21 faculty members and 42 non-faculty researchers.
@@ 2 Idea evaluation survey
For each wave, each participant completed a customized Google Forms survey lasting approximately 20 to 30 minutes. In each survey, participants evaluated seven ideas: six AI-generated ideas and one repeated human-written control idea (presented in random order). For each idea, participants rated eight scientific value dimensions: Originality, Paradigm relatedness, Acceptability, Implementability, Applicability, Effectiveness, Implicational explicitness, and Clarity, along with an Overall Quality score (0-10 scale). Detailed instructions and prompts are provided in SM 14.
@@ 3 Structured idea representations
When extracting ideas from existing papers, the pdf file was first converted to text using pymupdf4llm (?). Then the idea was extracted using an LLM model (gpt-4o-mini) with the prompt provided in SM 12. The temperature was set to 0.2 and top_p to 0.1. For both extracting and generating ideas, we used OpenAI’s structured output API to ensure the desired format is enforced.
@@ 4.1.2 Domain-compatibility pre-screening
Because not all cross-department pairings were equally promising, we performed a preliminary compatibility screen. For each department pair (e.g. [Biomedical Engineering, Sociology]), we sampled one conventional paper and ten candidate novel papers. For each of the ten pairs, we prompted the generator five times, yielding 50 ideas. We then scored these idea with our heuristic LLM evaluator (SM 4.2). The final compatibility of a pair of departments was calculated by the average LLM evaluator rating. Compatibility matrix is shown in Fig. S1
@@ 4.1.3 Model and prompting
We used OpenAI’s gpt-4o-2024-08-06 as the generator. For idea generation we kept temperature and top_p at their defaults (1.0) to encourage diversity. We used OpenAI’s structured output API to ensure that ideas were returned in a standardized structure SM 3. To avoid trivial re-statements of known work, the prompt explicitly instructed the model to be transformative, concrete, and not similar to existing works. The full generation prompt can be found in SM 13.
@@ 4.2 Evaluator design
Each generated idea was evaluated by an automated LLM-based evaluator that rated the same eight dimensions as human participants. To ground evaluations in existing literature, we implemented a retrieval-augmented pipeline. A gpt-4o-mini model generated five keyword summaries per idea, which were submitted to Semantic Scholar with five such queries. The retrieved papers’ titles, tldrs, and abstracts were saved into a pool of potentially relevant papers. We then filtered this pool of papers by their embedding similarity to the idea to be evaluated and kept those with a similarity higher than 0.88. We used OpenAI’s text-embedding-large model for this step. The threshold was decided after manual inspection for relevancy on a small set of examples.
The evaluation prompt (SM 14) used the same instructions and dimension anchors as those shown to human raters. For more consistent evaluation, we follow best practices in using LLMs as evaluators and compute a probability-weighted summation of the output token scores as the final score (?, ?, ?).
@@ 6 Predictors of rating change
Time gap. Since the time gap distribution was skewed (see SM 10), we took the log value of the days between responses.
@@ 7 Alignment of ideation system
9 Exploratory Study on Procedural Familiarity
To investigate whether procedural familiarity explains the drift, a separate $n=24$ non-faculty participants completed a fully-repeated condition. In this ablation group, all 7 AI plus human ideas in Wave 1 and Wave 2 are repeated. Across $24\times 7=168$ fully-repeated idea pairs, ratings showed a small and non-significant change between waves ($\Delta=+0.13$, 95% CI [0.24, 0.51][-0.24,\;0.51], $p=0.478$). This provides exploratory evidence that the survey procedure itself (repeated exposure to the same format) does not drive any systematic rating increase on its own.
@@ 14 Evaluation Instructions/Prompt
Table S1: Sensitivity analysis across analytical variants. Removed %: Percentage of the sample removed by the outlier screening method. ICC: test-retest reliability of Overall Quality for the repeated control idea — the primary, robust finding. Drift $\Delta$ / $p$: mean change and paired $t$-test for the control idea between waves; positive values indicate higher ratings in Wave 2. Raw AI $\Delta$: mean change in AI idea ratings before drift correction. DiD $p$: participant-level difference-in-differences $p$-value testing whether AI idea improvement exceeds concurrent drift in control idea ratings. ICC and DiD conclusions are consistent across all variants.
Version NN Removed % ICC Drift $\Delta$ Drift $p$ Raw AI $\Delta$ DiD $p$
no screening 63 0 $0.586$ +0.44+0.44 $0.121$ +0.29+0.29 $0.611$
IQR $3.0\times$ 63 0 $0.586$ +0.44+0.44 $0.121$ +0.29+0.29 $0.611$
IQR $1.5\times$ 62 1.6 $0.671$ +0.31+0.31 $0.226$ +0.29+0.29 $0.954$
IQR $1.0\times$ 60 4.8 $0.718$ +0.48+0.48 $0.036$ +0.34+0.34 $0.609$
$z\geq 3.0$ 62 1.6 $0.671$ +0.31+0.31 $0.226$ +0.29+0.29 $0.954$
$z\geq 2.5$ 62 1.6 $0.671$ +0.31+0.31 $0.226$ +0.29+0.29 $0.954$
$z\geq 2.0$ 59 6.3 $0.738$ +0.41+0.41 $0.064$ +0.36+0.36 $0.845$
MAD $\geq 3.5$ 62 1.6 $0.671$ +0.31+0.31 $0.226$ +0.29+0.29 $0.954$
@@
Originality 0.602 0.402 0.766
Paradigm relatedness 0.560 0.409 0.691
Acceptability 0.628 0.385 0.784
Implementability 0.524 0.240 0.739
Applicability 0.422 0.168 0.647
Effectiveness 0.543 0.328 0.707
Implicational explicitness 0.287 0.014 0.520
Clarity 0.285 0.017 0.538
@@
Originality 0.0 0.0 0.0 0.123 0.370 False
Paradigm relatedness 0.0 0.0 0.0 0.083 0.370 False
Acceptability 0.0 0.0 0.0 0.683 0.840 False
Implementability 0.0 0.0 0.0 0.840 0.840 False
Applicability 0.0 0.0 0.0 0.809 0.840 False
Effectiveness 0.0 0.0 0.0 0.139 0.370 False
Implicational explicitness 0.0 0.0 0.0 0.627 0.840 False
Clarity 0.0 0.0 0.0 0.383 0.767 False
Table S4: Mixed model on control Overall (drivers of change; standardized covariates)
@@
Intercept 6.045 0.411 14.723 0.000 5.240 6.849
Expertise 0.009 0.531 0.017 0.986 1.031-1.031 1.050
Wave 0.144 0.204 0.706 0.480 0.256-0.256 0.544
Position 0.054 0.132 0.410 0.682 0.205-0.205 0.313
Context 0.406-0.406 0.288 1.408-1.408 0.159 0.970-0.970 0.159
Time Gap 0.313-0.313 0.273 1.147-1.147 0.251 0.848-0.848 0.222
Time of Day 0.136 0.134 1.015 0.310 0.126-0.126 0.398
S-wave 1.839 0.209 8.778 0.000 1.428 2.249
@@
Figure S3: Full distribution of W1-W2 repeated idea drift values with IQR fences.